OpenStudy (anonymous):

what is the derivative of xsin(1/x)?

OpenStudy (anonymous):

product rule and chain rule for this one

OpenStudy (anonymous):

\[(fg)'-f'g+g'f\] \[\frac{d}{dx}[x\sin(\frac{1}{x})]=\sin(\frac{1}{x})+x\cos(\frac{1}{x})\times (-\frac{1}{x^2})\] clean up with some algebra